OpenStudy (anonymous):

Find the x- and y- intercepts of 5x+3y=15

OpenStudy (aravindg):

divide whole equation by 15

OpenStudy (aravindg):

you will get form x/a+y/b=1 where a is x intercept b is y intercept

OpenStudy (whpalmer4):

x-intercept is where y=f(x)=0. Solve 5x+3y=15 for y in terms of x, then set = 0 and solve for the x-intercept. y-intercept is where x = 0.

OpenStudy (whpalmer4):

AravindG's approach works well, too!

OpenStudy (whpalmer4):

My previous post ought to end with "y-intercept is value of y = f(0)"

OpenStudy (anonymous):

ok thank you